Advertisements
Advertisements
प्रश्न
Forty percent of business travellers carry a laptop. In a sample of 15 business travelers, what is the probability that atleast three of the travelers have a laptop?
Advertisements
उत्तर
Given n = 5
p = `40/10 = 2/5`
q = 1 – p = `1 - 2/5 = (5-2)/5 = 3/5`
The binomial distribution P(X = x) = `15"c"_x (2/5)^x (3/5)^(15 - x)`
P(at least three of the travelers have a laptop)
= P(X ≥ 3)
P(X ≥ 3) = 1 – P(X < 3)
= 1 – [P(X = 0) + P(X = 1) + P(X = 2)]
P(X < 3) = P(X = 0) + P(X = 1) + P(X = 2)
= `15"c"_0 (2/5)^0 (3/5)^(15 - 0) + 15"c"_1 (2/5)^1 (3/5)^(15 - 1) + 15"c"_2 (2/5)^2 (3/5)^(15 - 2)`
= `(1)(1)(3/5)^15 + 15(2/5)(3/5)^14 + 105(2/5)^2(3/5)^13`
= `13^15/5^15 +(15 xx 2 xx 3^14)/5^5 +(105 xx 4 xx 3^13)/5^15`
= `(3^13 [3^2 + (30 xx 3) + (105 xx 4 xx 9)])/5^15`
= `(1.593 xx 10^6 xx (9 + 90 + 3780))/(3.055 xx 10^10)`
= `(1.593 xx 3879)/(3.055 xx 10^4)`
= `6179.247/30550`
= 0.20226
∴ P(X ≥ 3) = `1 - 020226`
= 0.79774
APPEARS IN
संबंधित प्रश्न
In a particular university 40% of the students are having newspaper reading habit. Nine university students are selected to find their views on reading habit. Find the probability that atleast two-third have newspaper reading habit
Write any 2 examples for Poisson distribution
Mention the properties of poisson distribution
Write down any five chief characteristics of Normal probability curve
In a test on 2,000 electric bulbs, it was found that bulbs of a particular make, was normally distributed with an average life of 2,040 hours and standard deviation of 60 hours. Estimate the number of bulbs likely to burn for less than 1,950 hours
Choose the correct alternative:
The parameters of the normal distribution f(x) = `(1/sqrt(72pi))"e"^(-(x - 10)^2)/72 - oo < x < oo`
Choose the correct alternative:
Monthly expenditure on their credit cards, by credit cardholders from a certain bank, follows a normal distribution with a mean of ₹ 1,295.00 and a standard deviation of ₹ 750.00. What proportion of credit cardholders spend more than ₹ 1,500.00 on their credit cards per month?
A manufacturer of metal pistons finds that on the average, 12% of his pistons are rejected because they are either oversize or undersize. What is the probability that a batch of 10 pistons will contain at least 2 rejects?
Hospital records show that of patients suffering from a certain disease 75% die of it. What is the probability that of 6 randomly selected patients, 4 will recover?
Entry to a certain University is determined by a national test. The scores on this test are normally distributed with a mean of 500 and a standard deviation of 100. Raghul wants to be admitted to this university and he knows that he must score better than at least 70% of the students who took the test. Raghul takes the test and scores 585. Will he be admitted to this university?
